two silver alloys contain 40% and 30% silver. how many kilograms of each alloy should be mixed to obtain 20 kilograms of 36% SILVER ALLOY?
Answer by mananth(12270)

percent ---------------- quantity
Alloy I 40 ---------------- x Kg
Alloy II 30 ------ 20 - x Kg
Mixture 36 ---------------- 20

40x+30 (20-x)= 36*20
4000x+ 3000(20 -x)= 72000
4000x+ 60000-3000x= 72000
4000x- 3000x=72000-60000
1000 x = 12000
/ 1000
x = 12 Kg 40.00% Alloy I
8 Kg 30.00% Alloy II

m.ananth@hotmail.ca